OpenStudy (anonymous):

Find the sum of the summation of 3i plus 2, from i equals 5 to 13.

OpenStudy (anonymous):

Find the sum of \[\sum_{i=5}^{13}3i+2\]

OpenStudy (anonymous):

you cant put this in a calculator

OpenStudy (anonymous):

For some reason I keep getting it wrong can you tell me what you got

OpenStudy (anonymous):

ok I will hold on

OpenStudy (anonymous):

what did you get though

OpenStudy (anonymous):

@ihateschool18 :very simple: \[\sum_{i=5}^{13} = \sum 17+\sum 20+\sum 23+\sum 26+\sum 29+\sum 32+\sum 35+\sum 38+\sum 41\]

OpenStudy (anonymous):

\[=261\]

OpenStudy (anonymous):

Thank you guys very much.